# Real-World Data and Calibrated Simulation Suite for Offline Training of Reinforcement Learning Agents to Optimize Energy and Emission in Buildings for Environmental Sustainability

### v1

@@ Abstract
Commercial office buildings contribute 17 percent of Carbon Emissions in the US, according to the US Energy Information Administration (EIA), and improving their efficiency will reduce their environmental burden and operating cost. A major contributor of energy consumption in these buildings are the Heating, Ventilation, and Air Conditioning (HVAC) devices. HVAC devices form a complex and interconnected thermodynamic system with the building and outside weather conditions, and current setpoint control policies are not fully optimized for minimizing energy use and carbon emission. Given a suitable training environment, a Reinforcement Learning (RL) agent is able to improve upon these policies, but training such a model, especially in a way that scales to thousands of buildings, presents many practical challenges. Most existing work on applying RL to this important task either makes use of proprietary data, or focuses on expensive and proprietary simulations that may not be grounded in the real world. We present the Smart Buildings Control Suite, the first open source interactive HVAC control dataset extracted from live sensor measurements of devices in real office buildings. The dataset consists of two components: six years of real-world historical data from three buildings, for offline RL, and a lightweight interactive simulator for each of these buildings, calibrated using the historical data, for online and model-based RL. For ease of use, our RL environments are all compatible with the OpenAI gym environment standard. We also demonstrate a novel method of calibrating the simulator, as well as baseline results on training an RL agent on the simulator, predicting real-world data, and training an RL agent directly from data. We believe this benchmark will accelerate progress and collaboration on building optimization and environmental sustainability research.
@@ 1 Introduction
Energy optimization and management in commercial buildings is a very important problem, whose importance is only growing with time. Buildings account for 37% of all US carbon emissions, with commercial buildings alone taking up a staggering 17% in 2023 (EIA, ). Reducing those emissions by even a small percentage can have a significant effect. In climates that are either very hot or very cold, energy consumption is much higher, and there is even more room to have a major impact. We believe this problem is one of the most important avenues for climate sustainability research, where even a small improvement over baseline policies can drastically reduce our carbon footprint.
@@
In particular, HVAC systems account for 40-60% of energy use in buildings (Pérez-Lombard et al., 2008), and roughly 15% of the world’s total energy consumption (Asim et al., 2022). Most office buildings are equipped with advanced HVAC devices, like Variable Air Volume (VAV) devices, Hot Water Systems, Air Conditioners and Air Handlers that are configured and tuned by the engineers, manufacturers, installers, and operators to run efficiently with the device’s local control loops (McQuiston et al., 2023). However, integrating multiple HVAC devices from diverse vendors into a building “system” requires technicians to program fixed operating conditions for these units, which may not be optimal for every building and every potential weather condition. Existing setpoint control policies are not optimal under all conditions, and the possibility exists that a machine learning model may be trained to continuously tune a small number of setpoints to achieve greater energy efficiency and reduced carbon emission.
@@
Optimizing HVAC control has been an active research area for decades, and yet while AI has begun to revolutionize many industries, to date almost all HVAC systems remain the same as they were 30 years ago: despite all the literature on the topic, there is not a single solution that has been widely adopted in the real world.
@@
One of the most significant factors limiting progress is the lack of a reliable public benchmark to test solutions against. Current work generally makes use of proprietary data and expensive (often also proprietary) simulations. This limits participation to those with exclusive access, and makes most claims difficult to verify and compare. A strong public dataset would facilitate collaborations between institutions, standardize research efforts, and allow for wider participation. Historically, much of progress in AI has been driven by easily accessible public benchmarks, from the ImageNet Challenge in Vision (Russakovsky et al., 2015), to the Atari57 suite in RL (Badia et al., 2020), and the GLUE Benchmark in language (Wang et al., 2018). A similar benchmark in HVAC control may help accelerate progress and finally lead to adoption of solutions in the real world.
@@
We present The Smart Buildings Control Suite, a high quality, fully accessible, building control benchmark. The benchmark consists of two components:
@@
  • Real-world historical HVAC data, collected from three buildings over a six year period.
@@
  • A highly customizable and scalable HVAC and building simulator, with configurations corresponding to each of the above buildings
Our contributions include one of the first public real-world HVAC datasets, a highly customizable and scalable HVAC and building simulator, a rapid configuration method to customize the simulator to a particular building, a calibration method to improve this fidelity using real-world data, and an evaluation method to measure the simulator fidelity. The dataset contains information from three buildings in California, the largest of which is three stories and 118,086 ft2. Using data we obtained from each building, we calibrate our simulator, and demonstrate using our evaluation pipeline that this significantly improves its fidelity to the real building. We provide pre-calibrated simulators for all of our buildings, as well as code to both reproduce the calibration procedure, and to calibrate the simulator to new scenarios. While our suite focuses on three buildings, our simulator is easily adaptable, allowing for the development of general purpose solutions that can be applied to any building. All the data and simulator code is open source and compatible with the OpenAI gym environment standard(Brockman et al., 2016), and data is available on the popular TensorFlow Datasets platform (TFDS, ) under the Creative Commons License.
We first give an overview of the problem and related work, and then present the structure of the data. Next we introduce the simulator, and discuss our configuration, calibration, and evaluation techniques. After that, we run through an example of the process of calibrating the simulator to real data, and finally we demonstrate success on three key benchmark tasks: training an RL agent on the calibrated simulator environment using Soft Actor Critic(Haarnoja et al., 2018), training a regression model to predict the real world dynamics, and training a Soft Actor Critic agent from the real world data via the regression model.
2 Optimizing Energy and Emission in Office Buildings with RL
@@
In this section we frame energy optimization in office buildings as an RL problem. We define the state of the office building $S_{t}$ at time $t$ as a fixed length vector of measurements from sensors on the building’s devices, such as a specific VAV’s zone air temperature, gas meter’s flow rate, etc. The action on the building $A_{t}$ is a fixed-length vector of device setpoints selected by the agent at time $t$, such as the boiler supply water temperature setpoint, etc.
More generally, RL is a branch of machine learning that attempts to train an agent to choose the best actions to maximize some long-term, cumulative reward (Sutton & Barto, 2018). The agent observes the state $S_{t}$ from the environment at time $t$, then chooses action $A_{t}$. The environment responds by transitioning to the next state $S_{t+1}$ and returns a reward (or penalty) after the action, $R_{t+1}$. Over time, the agent will explore the action space and learn to maximize the reward over the long term for each given state. A discount factor $\gamma$ reduces the value of future rewards amplifying the value of the near-term reward. When this cycle is repeated over multiple episodes, the agent converges on a state-action policy that maximizes the long-term reward.
@@
This sequence is often formalized as the Markov Decision Process (MDP) (Garcia & Rachelson, 2013), described by the tuple $(S,A,p,R)$ where the state space is continuous (e.g., temperatures, flow rates, etc.) and the action space is continuous (e.g., setpoint temperatures) and the transition probability $p:S\times S\times A\rightarrow[0,1]$ represents the probability density of the next state $S_{t+1}$ from taking action $A_{t}$ on the current state $S_{t}$. The reward function $R:S\times A\rightarrow[R_{min},R_{max}]$ emits a single scalar value at each time $t$. The agent is acting under a policy $\pi_{\theta}(A_{t}|S_{t})$ parameterized by $\theta$ that represents the probability of taking action $A_{t}$ from state $S_{t}$. The goal of an RL agent is to find the policy that maximizes the expected long-term cumulative, discounted reward. The set of parameters $\bm{\theta}^{*}$ of the optimal policy can be expressed as:
$\bm{\theta}^{*}=\arg\max_{\theta}\mathbb{E}_{\tau\sim\pi_{\theta}(\tau)}\left[\sum_{t}\gamma^{t}R(S_{t},A_{t})\right]$
where $\theta$ is the current policy parameter, and $\tau$ is a trajectory of states, actions, and rewards over sequential time steps $t$. In order to converge to the optimal policy, the agent requires many training iterations to explore the policy space, making online training directly on the real-world building from scratch inefficient, dangerous, impracticable, and likely impossible. Therefore, it is necessary to enable offline learning, where the agent can train in an efficient sandbox environment that adequately emulates the dynamics of the building before being deployed to the real world.
@@
Reward Function RL generally requires a single scalar reward signal, $R_{t}(S_{t},A_{t})$ that indicates the quality of taking action $A_{t}$ in state $S_{t}$. We thus define a custom feedback signal, $R_{3C}$, as a weighted sum of negative cost functions for carbon emission, energy cost, and comfort levels within the building, which we dubbed the 3C Reward. It is governed by the following equation:
@@
where $C_{1}$ represents normalized comfort conditions, $C_{2}$ normalized energy cost and $C_{3}$ normalized carbon emission. Constants $u$, $v$, $w$ represent operator preferences, allowing them to weight the relative importance of cost, comfort and carbon consumption. $R_{3C}=0$ when no energy is consumed, no carbon is emitted, and all occupied zones are in setpoint bounds, and negative otherwise. For more details, and equations governing how we normalize and measure these quantities, see Appendix A.
3 Related Works
Considerable attention has been paid to HVAC control (Fong et al., 2006) in recent years (Kim et al., 2022), and while alternative approaches exist, such as model predictive control (Taheri et al., 2022), a growing portion of the literature has considered how RL and its various associated algorithms can be leveraged (Yu et al., 2021; Mason & Grijalva, 2019; Yu et al., 2020; Gao & Wang, 2023; Wang et al., 2023; Vázquez-Canteli & Nagy, 2019; Zhang et al., 2019b; Fang et al., 2022; Zhang et al., 2019b). As mentioned above, a central requirement in RL is the offline environment that trains the RL agent. Several methods have been proposed, largely falling under three broad categories.
Data-driven Emulators Some works attempt to learn a dynamics as a multivariate regression model from real-world data (Zou et al., 2020; Zhang et al., 2019a), often using recurrent neural network architecture, such as Long Short-Term Memory (LSTM) (Velswamy et al., 2017; Sendra-Arranz & Gutiérrez, 2020; Zhuang et al., 2023). The difficulty here is that data-driven models often do not generalize well to circumstances outside the training distribution, especially since they are not physics based.
Offline RL The second approach is to train the agent directly from the historical real-world data, without ever producing an interactive environment (Chen et al., 2020; 2023; Blad et al., 2022). While the real-world data is obviously of high accuracy and quality, this presents a major challenge, since the agent cannot take actions in the real world and interact with any form of an environment. This inability to explore severely limits its ability to improve over the baseline policy producing the real-world data (Levine et al., 2020). Furthermore, prior to our work, there are few public datasets available.
Physics-based Simulation HVAC system simulation has long been studied (Trčka & Hensen, 2010; Riederer, 2005; Park et al., 1985; Trčka et al., 2009; Husaunndee et al., 1997; Trcka et al., 2007; Blonsky et al., 2021). EnergyPlus (Crawley et al., 2001), a high-fidelity simulator developed by the Department of Energy, is commonly used (Wei et al., 2017; Azuatalam et al., 2020; Zhao et al., 2015; Wani et al., 2019; Basarkar, 2011), but suffers from scalability and configuration challenges.
To overcome the limitations of each of the above three methods, some work has proposed a hybrid approach (Zhao et al., 2021; Balali et al., 2023; Goldfeder & Sipple, 2023; Zhang et al., 2023; Klanatsky et al., 2023; Drgoňa et al., 2021), and indeed this is the category our work falls under. What is unique about our approach is the use of a physics based simulator that achieves an ideal balance between speed of configuration, and fidelity to the real world. Our simulator is lightweight enough to be configured to an arbitrary building in a matter of hours, and using our calibration process based on real-world data, accurate enough to train an effective control agent off-line. This allows our solution to be highly scalable, like the first two approaches, but still rooted in physics, and demonstrably calibrated, like the third approach.
Various works have also discussed how exactly to apply RL to an HVAC environment, such as what sort of agent to train. Inspired by prior effective use of Soft Actor Critic (SAC) on related problems (Kathirgamanathan et al., 2021; Coraci et al., 2021; Campos et al., 2022; Biemann et al., 2021), we chose to demo our environment using a SAC agent.
Prior Datasets While many building datasets exist (Ye et al., 2019), most either have a different focus (Sachs et al., 2012; Urban et al., 2015; Kriechbaumer & Jacobsen, 2018; Granderson et al., 2023), do not contain sufficient HVAC information (Miller et al., 2020; Mathew et al., 2015; Rashid et al., 2019; Jazizadeh et al., 2018; Sartori et al., 2023), are focused on residential buildings (Murray et al., 2017; Barker et al., 2012; Meinrenken et al., 2020) or non-standard buildings (Pettit et al., 2014; Biswas & Chandan., 2022), or are simulated (Field et al., 2010; Bakker et al., 2022). Even the few datasets directly relevant (Luo et al., 2022; Heer et al., 2024) are non-interactive. As far as we are aware, we present the first HVAC control benchmark that has high quality real-world data with computationally cheap simulations of the same buildings, allowing for both real-world grounding and interactive control experiments.
@@ 3 Related Works
Figure 1: Example Visualization of an Environment. Blue represents colder temperatures, red warmer. Blue and red dots inside the building indicate diffusers that are dispensing cold and warm air respectively.
4 The Dataset Structure

### v2

# The Smart Buildings Control Suite: A Diverse Open Source Benchmark to Evaluate and Scale HVAC Control Policies for Sustainability

@@ Abstract
Commercial buildings account for 17% of U.S. carbon emissions, with roughly half of that from Heating, Ventilation, and Air Conditioning (HVAC). HVAC devices form a complex thermodynamic system, and while Model Predictive Control and Reinforcement Learning have been used to optimize control policies, scaling to thousands of buildings remains a significant unsolved challenge. Most current algorithms are over-optimized for specific buildings and rely on proprietary data or hard-to-configure simulations. We present the Smart Buildings Control Suite, the first open source interactive HVAC control benchmark with a focus on solutions that scale. It consists of 3 components: real-world telemetric data extracted from 11 buildings over 6 years, a lightweight data-driven simulator for each building, and a modular Physically Informed Neural Network (PINN) building model as a simulator alternative. The buildings span a variety of climates, management systems, and sizes, and both the simulator and PINN easily scale to new buildings, ensuring solutions using this benchmark are robust to these factors and only reliant on fully scalable building models. This represents a major step towards scaling HVAC optimization from the lab to buildings everywhere. To facilitate use, our benchmark is compatible with the Gym standard, and our data is part of TensorFlow Datasets.
Machine Learning, ICML
@@ 1 Introduction
Energy optimization and management in commercial buildings is a very important problem, whose importance is only growing with time. Buildings account for 37% of all US carbon emissions, with commercial buildings alone taking up a staggering 17% in 2023 (EIA, ). Reducing those emissions by even a small percentage can have a significant effect, especially in more extreme climates. We believe this problem is one of the most important avenues for climate sustainability research, where even a small improvement over baseline policies can drastically reduce our carbon footprint.
@@
In particular, HVAC systems account for 40-60% of energy use in buildings (Pérez-Lombard et al., 2008) and roughly 15% of the world’s total energy consumption (Asim et al., 2022). Most office buildings are equipped with advanced HVAC devices, like Variable Air Volumes (VAVs), Hot Water Systems, Air Conditioners (ACs) and Air Handlers (AHUs) that are configured and tuned by the engineers, manufacturers, installers, and operators to run efficiently with the device’s local control loops (McQuiston et al., 2023). However, integrating multiple HVAC devices from diverse vendors into a building “system” requires technicians to program fixed operating conditions for these units, which may not always be optimal, and thus an ML model can be trained to continuously tune a small number of setpoints to achieve greater energy efficiency and reduced carbon emission.
@@
Optimizing HVAC control has been an active research area for decades, and yet while AI has begun to revolutionize many industries, to date almost all HVAC systems remain the same as they were 30 years ago: despite all the literature on the topic, there is not a single solution that has been widely adopted in the real world. One of the most significant factors limiting progress is the absence of a reliable public benchmark to evaluate solutions. Current efforts often rely on proprietary data and expensive closed-source simulations. This restricts participation to those with exclusive access and makes it challenging to verify and compare results. Another major challenge is the lack of scalability across different buildings. Most solutions are tailored to specific buildings and fail to generalize due to two primary reasons:
  1. 1.
    Complex building models: Many approaches rely on high-fidelity simulators or models that are difficult and time-consuming to configure for arbitrary buildings.
  2. 2.
    System-specific dependencies: Solutions are often tied to a particular building management system, HVAC system design, or data ontology, making them brittle and unable to adapt to the variability of other systems.
@@
A strong and diverse public benchmark would facilitate collaborations between institutions, standardize research efforts, allow for wider participation, and sharpen the focus on scalable solutions. Historically, much of the progress in AI has been driven by easily accessible public benchmarks, from the ImageNet Challenge in Vision (Russakovsky et al., 2015), to the Atari57 suite in RL (Badia et al., 2020), and the GLUE Benchmark in NLP (Wang et al., 2018). A similar benchmark in HVAC control may help accelerate progress and finally lead to the adoption of solutions in the real world.
@@
We present The Smart Buildings Control Suite, a diverse, high quality, fully accessible, building control benchmark; the first of its kind. It consists of the following components:
  1. 1.
@@
    Real-world historical HVAC data, collected from 11 buildings spanning 3 building management systems from across North America, over a 6-year period.
  2. 2.
@@
    A highly customizable and scalable HVAC and building simulator that can be calibrated with real data, with configurations corresponding to each of the above buildings, as well as a pipeline for easily onboarding and calibrating new buildings.
  3. 3.
    A modularized neural network architecture incorporating physical priors for building energy modeling, as a fully data-driven, physically informed neural network (PINN) simulator alternative, as well as models trained to emulate each of the above buildings, and instructions for training a model on a new building.
  4. 4.
    A focus on ease of use. This includes full compatibility with the OpenAI Gym standard(Brockman et al., 2016) so that a model can be trained either on offline real data or interactively from the simulator or PINN backend. Our data is also available on the popular TensorFlow Datasets platform (TFDS, ), and our code is open source.
This benchmark is one of the first public real-world HVAC datasets, and certainly the first with a clear focus on scaling to varied buildings. Our data comes from diverse sources and climates, and we present two novel building modeling methods, both data-driven and directly designed to easily scale to new and diverse building types. Thus, solutions developed on this benchmark should be robust, and not reliant on building models that do not scale. We first discuss related work, and give an overview of the problem setup and various optimization approaches and considerations. Then, we present the real-world dataset. Next we introduce the simulator, discuss our configuration, calibration, and evaluation techniques, and run through an example of the simulator calibration process. After, we describe the PINN model and demonstrate its predictive performance as compared with an LSTM. Finally, we present an example of learning a control policy using our benchmark that improves over the baseline.
2 Related Work
Considerable attention has been paid to HVAC control (Fong et al., 2006) in recent years (Kim et al., 2022), and while alternative approaches exist, such as MPC (Taheri et al., 2022), a growing portion of the literature has considered how RL and can be leveraged (Yu et al., 2021; Mason & Grijalva, 2019; Yu et al., 2020; Gao & Wang, 2023; Wang et al., 2023; Vázquez-Canteli & Nagy, 2019; Zhang et al., 2019b; Fang et al., 2022; Zhang et al., 2019b; Goldfeder & Sipple, 2024). As mentioned above, a central requirement is the offline environment that trains the RL agent. Several methods have been proposed, largely falling under three broad categories. Our benchmark correspondingly has three parts, representing the state of the art of each.
Offline RL on Real Data The first approach is to train the agent directly from the historical real-world data, without ever producing an interactive environment (Chen et al., 2020, 2023a; Blad et al., 2022). While the real-world data is obviously of high accuracy and quality, this presents a major challenge, since the agent cannot take actions in the real world and interact with any form of an environment. This inability to explore severely limits its ability to improve over the baseline policy producing the real-world data (Levine et al., 2020). Furthermore, prior to our work, there are few public datasets available. Our dataset, with its diverse buildings and long duration, should allow for rapid development of offline RL agents in a way not previously possible.
Data-driven Emulators Some works attempt to learn dynamics as a multivariate regression model from real-world data (Zou et al., 2020; Zhang et al., 2019a), often using recurrent neural network architecture, such as Long Short-Term Memory (LSTM) (Velswamy et al., 2017; Sendra-Arranz & Gutiérrez, 2020; Zhuang et al., 2023). The difficulty here is that data-driven models often do not generalize well to circumstances outside the training distribution, especially since they are not physics-based. To overcome this limitation, Physically Informed Neural Networks (PINNS) incorporate physical priors to enforce reasonable behavior in out-of-distribution scenarios (Djeumou et al., 2022; Wang & Dong, 2023; Chen et al., 2023b; Gokhale et al., 2022; Jiang & Dong, 2024), and our PINN model builds off the state of the art in this area.
Physics-based Simulation HVAC system simulation has long been studied (Trčka & Hensen, 2010; Riederer, 2005; Park et al., 1985; Trčka et al., 2009; Husaunndee et al., 1997; Trcka et al., 2007; Blonsky et al., 2021). EnergyPlus (Crawley et al., 2001), a high-fidelity simulator developed by the Department of Energy, is commonly used (Wei et al., 2017; Azuatalam et al., 2020; Zhao et al., 2015; Wani et al., 2019; Basarkar, 2011), but suffers from scalability and configuration challenges. To overcome the limitations of each of the above methods, some work has proposed a hybrid approach (Zhao et al., 2021; Balali et al., 2023; Goldfeder & Sipple, 2023; Zhang et al., 2023; Klanatsky et al., 2023; Drgoňa et al., 2021), and indeed this is the category our simulator falls under. What is unique about our approach is the use of a physics-based simulator that achieves an ideal balance between speed of configuration, and fidelity to the real world. Our simulator is lightweight enough to be configured rapidly to an arbitrary building, easy to calibrate with data, and accurate enough to train an effective controller.
Prior Datasets While many building datasets exist (Ye et al., 2019), most either have a different focus (Sachs et al., 2012; Urban et al., 2015; Kriechbaumer & Jacobsen, 2018; Granderson et al., 2023), do not contain sufficient HVAC information (Miller et al., 2020; Mathew et al., 2015; Rashid et al., 2019; Jazizadeh et al., 2018; Sartori et al., 2023), are focused on residential buildings (Murray et al., 2017; Barker et al., 2012; Meinrenken et al., 2020) or non-standard buildings (Pettit et al., 2014; Biswas & Chandan., 2022), or are simulated (Field et al., 2010; Bakker et al., 2022). Even the few datasets directly relevant (Luo et al., 2022; Heer et al., 2024) are non-interactive and come from a single building management system. We present the first HVAC control benchmark that has high quality real-world data from diverse sources, along with computationally cheap data-driven simulation and PINN building models, allowing for both real-world grounding and interactive control experiments.
3 Optimizing Energy and Emission in Commercial Buildings
3.1 Problem Formulation
@@
In this section, we frame energy optimization in buildings as a Markov Decision Process (MDP)(Garcia & Rachelson, 2013). We define the state of the building $S_{t}$ at time $t$ as a fixed length vector of measurements from sensors on the building’s devices, such as a specific VAV’s zone air temperature, gas meter’s flow rate, etc. The action on the building $A_{t}$ is a fixed-length vector of device setpoints selected by the controller at time $t$, such as the boiler supply water temperature setpoint, etc. The controller observes the state $S_{t}$ from the environment at time $t$, then chooses action $A_{t}$. The environment responds by transitioning to the next state $S_{t+1}$ and returns a reward after the action, $R_{t+1}$.
@@
Thus the MDP is formally described by the tuple $(S,A,p,R)$ where the state space is continuous (e.g., temperatures, flow rates, etc.) and the action space is continuous (e.g., setpoint temperatures) and the transition probability $p:S\times S\times A\rightarrow[0,1]$ represents the probability density of the next state $S_{t+1}$ from taking action $A_{t}$ on the current state $S_{t}$. The reward function $R:S\times A\rightarrow[R_{min},R_{max}]$ emits a scalar at each time $t$. The controller is acting under policy $\pi_{\theta}(A_{t}|S_{t})$ parameterized by $\theta$ that represents the probability of taking action $A_{t}$ from state $S_{t}$. The MDP formalism is broad and allows for many optimization strategies, such as Reinforcement Learning (RL), heuristics, and Model Predictive Control (MPC). For an overview of approaches to learning an optimal policy, see Appendix A.
3.2 Reward Function
@@
The MDP formalism generally requires a single scalar reward signal, $R_{t}(S_{t},A_{t})$ that indicates the quality of taking action $A_{t}$ in state $S_{t}$. Since this is a multi-objective optimization problem, we define a custom feedback signal, $R_{3C}$, as a weighted sum of negative cost functions for carbon emission, energy cost, and comfort levels within the building, which we call the 3C Reward. It is governed by the following equation:
@@
where $C_{1}$ represents normalized comfort conditions, $C_{2}$ normalized energy cost and $C_{3}$ normalized carbon emission. Constants $u$, $v$, $w$ represent operator preferences, allowing them to weigh the relative importance of cost, comfort, and carbon consumption. $R_{3C}=0$ represents the best-case scenario: no energy is consumed, no carbon is emitted, and all occupied zones are in setpoint bounds. For details and equations governing how we measure and normalize these quantities, see Appendix B.
Indoor air quality is an important dimension not represented in the 3C Reward. An additional reward term can be optionally added to account for it during training, and an air quality constraint can also be used at deployment time. Appendix B further quantifies air quality as a potential reward term or constraint, informed by the ANSI/ASHRAE Standard 62.1 standard for indoor air quality (ASHRAE, 2022).
@@
Figure 1: Visualization of an Environment. Colder temperatures are blue; warmer ones are red. Blue and red dots inside the building indicate diffusers dispensing cold and warm air respectively.
4 The Smart Buildings Dataset